i have two solution sets from two linear equations
equation one:
y=4(x)-1
The solution set is:
(-2, -9)
(0, -1)
(2, 7)
(5, 19)
Equation 2:
y=3(x)+5
The solution set is:
(-2, 9)
(0,15)
(2, 21)
(5, 30)
As you can see I have already found the solution set; my problem is that once I graph it I can not get them to cross to show the solution point. Can you please help me?

Found 2 solutions by RAY100, solver91311:
Answer by RAY100(1637)

Good work,,,You understand how graphing works,
.
The first set looks good
.
looking at the second set,,,y = 3x +5
for x=0, y =0 +5 = 5
for x=2, y= 3(2) +5 = 11
for x=5, y=3(5) +5 = 20
for x=-2, y=3(-2) +5 = -1
.
if we look at the two eqns
1) y=4x-1
2) y=3x+5
.
4x-1 =3x+5
x=6
substituting in (1)
y = 4(6) -1 = 23
.
solution ( intersection) at (6,23)
.
checking
substitute in both eqns
1) 23 = 4(6)-1,,,,ok
2) 23 = 3(6) +5,,,,ok

Answer by solver91311(24713)


Since y is equal to two different things, why not set them equal to each other?



Add -3x to both sides, add 1 to both sides:



Then



Point of intersection is (6, 23)




John